The standard mass of an unknown substance is 15.6g. Which of the following set has the least accurate value? Show solutions



Trial A B C


1 15.6 15.8 15.6


2 15.4 14.6 15.5


3 15.6 14.7 15.8


1
Expert's answer
2021-09-29T17:51:19-0400

"(15.6+15.8+15.6)\/3=15.7"


"(15.4+14.6+15.5)\/3=15.2"


"(15.6+14.7+15.8)\/3=15.4"


The least accurate value 15.6